OpenStudy (anonymous):

got one more...6x^3+4x^2+16, when x=2

OpenStudy (anonymous):

the value of the function, when x=2, is 80

OpenStudy (anonymous):

6*(2)^3+4*(2)^2+16=80... simple calculation..

OpenStudy (anonymous):

it's like $$ 6 \times 8 + 4 \times 4 +16 = 48+32 = 80 $$

OpenStudy (anonymous):

true..tried to make it harder than it was

OpenStudy (anonymous):

Do you have anything else ?